Fermi National Laboratory
 

Operation Summary for August 26, 2011 - August 29, 2011

 

The Integrated Luminosity for the period from 8/22/11 to 8/29/11 was 26.25 pb-1 (inverse picobarns). NuMI reported receiving 5.55E18 protons on target during this same period.

Friday August 26

The day shift began with Operations monitoring the Tevatron (TEV) store 9032 (58.36E30) and the Recycler stash (159.5E10), with the Antiproton Source (Pbar) stacking (192.1E10), with the Main Injector (MI) operating normally, but with the Pelletron off for repairs, with Linac and Booster operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, the Neutron Therapy Facility (NTF) reported at 8:30 AM that they had begun treating the first of two patients for the day. At 8:56 AM, a Linac RF station (LRF3) tripped off for the first of many times in an hour; Linac technicians investigated and managed to keep the station running for NTF, but not for HEP. NTF completed treating their last patient and Operations terminated store 9032 around 11:36 AM. Operations loaded a 3x3 study store into the TEV around 12:21 PM. Operators prepared Booster for an access at 12:37 PM for RF work. Linac personnel reported at 2:45 PM that they had completed their work on LRF3. Operations began sending beam to MTA at 3:18 PM for studies.

During the evening shift, the Booster RF work ended at 4:43 PM; Operations resumed sending beam to NuMI and Pbar. At 5:04 PM, operators reported that they could not keep the MiniBooNE horn on; they called in horn personnel. Operations established beam to the Meson FTBF T-1015 experiment at 5:24 PM. Operations disabled beam to MTA at 5:38 PM. Operations terminated the 3x3 study store at 5:38 PM. Operations put the TEV through a wet squeeze at 6:11 PM. Operations put the TEV into shot setup at 6:36 PM. MiniBooNE horn personnel reported at 6:58 PM that they had found the signal amplitude for Cell 11 was down by 8%; they reconnected cables and the loss went away. Operations established store 9035 at 7:54 PM with an average initial luminosity (AIL) of 115.01E30. Operations disabled beam to T-1015 at 8:12 PM.

Saturday August 27

The midnight shift began with the Pelletron off, with Operations monitoring store 9035 (61.68E30) and the stash (159.9E10), with Pbar stacking (119.3E10), and with the NuMI and MiniBooNE taking beam.

During the midnight shift, at 12:09 AM, LRF3 tripped off for the first of three times in four minutes; Operations put Pbar in standby while they brought LRF3 up slowly. Pelletron personnel reported at 12:10 AM that they had completed the SF6 gas transfer into the tank. Operators had the MI-31/Pelletron enclosure searched and secured by 1:34 AM. Pelletron personnel had the Pelletron online by 2:45 AM. Operations terminated store 9035 at 3:32 AM, put the TEV into shot setup, and established store 9036 at 4:49 AM with an AIL of 115.2E30. At 6 AM, operators reported that a transfer of antiprotons did not make it from the Pbar Accumulator to the Recycler due to a kicker (R:KPS3A) trip; operators reset the kicker.

During the day shift, the Pelletron recovery continued. At 9:01 AM, the Pelletron suffered a full discharge. Pelletron personnel reported at 9:47 AM that the level three electronics had failed again; operators began preparing the Pelletron for access and called in a crew to remove the SF6 insulating gas from the tank. Operations terminated store 9036 at 11:35 AM, put the TEV into shot setup, and established store 9037 at 1 PM with an AIL of 102.43E30. At 2:11 PM, operators noticed a Booster beam efficiency problem and began an injection tune. At 3:07 PM, operators restored a backup file, but this did not help Booster beam efficiency, so operators ran another injection tune, which was also of no help.

During the evening shift, at 7:16 PM, the Pbar LCW system alarmed; operators toured the service buildings, but found no indication of a leak. Operations disabled the beam to T-1015 at 9 PM. Operations terminated store 9037 at 9:03 PM and put the TEV into shot setup. Pelletron work began around 9:30 PM. Operators reported at 9:37 PM that they had found and replaced a bad Booster quadrupole magnet power supply control card, which caused the Booster beam efficiency problems. Operations established store 9038 at 10:19 PM with an AIL of 97.7E30. The MI RF system tripped off on a LCW pressure indication; operators investigated, contacted a RF person, and then changed a nitrogen bottle, which controls the LCW pressure.

Sunday August 28

The midnight shift began with the Pelletron off for work, with Operations monitoring store 9038 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, operators reported at 1:32 AM that the Pelletron work had finished; operators called in personnel to fill the tank with SF6 gas. At 2:37 AM, BRF15 tripped off on a ground fault indication and would not reset; operators bypassed the station, but eventually brought it back online. Operations terminated store 9038 at 7:07 AM and put the TEV into shot setup. At 7:52 AM, operators reported that during a transfer to the TEV, the second bunch of antiprotons made it through MI, but not into the TEV; operators reported that the TEV either kicked or cogged at the wrong spot. Operators took the TEV out of shot setup and put it through a dry squeeze at 7:59 AM.

During the day shift, a TEV person reported at 8:12 AM that the TEV sector E48 kicker had prefired; he was going to raise the gas pressure for the kicker. Operations put the TEV into shot setup at 9:20 AM and established store 9040 at 10:17 AM with an AIL of 92.53E30. Operations established beam to T-1015 at 10:39 AM. At 11:32 AM, LRF2 tripped off for the first of two times in 50 minutes; it reset. At 1:30 PM, LRF3 tripped off for the first of 6 times in 39 minutes; it reset.

During the evening shift, Operations terminated store 9040 at 5:34 PM, put the TEV into shot setup, and established store 9041 at 6:42 PM with an AIL of 116.53E30. Pelletron personnel put the Pelletron into conditioning mode at 7:10 PM. Operations disabled beam to T-1015 at 9:02 PM. A Pelletron person declared the Pelletron was ready for operation.

Monday August 29

The midnight shift began with the Pelletron operating, with Operations monitoring store 9041 and the stash, with Pbar stacking, and with the NuMI and MiniBooNE taking beam.

During the midnight shift, at 12:35 AM, LRF3 tripped off for the first of three times in six minutes; the station would not reset after the last trip. Operators called a Linac person; she had the operators change the gradients for LRF2 and LRF4. Operations terminated store 9041 at 4:27 PM, put the TEV into shot setup, and established store 9042 at 5:42 AM with an AIL of 262.4E30.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster studies, MI coalescing studies, and a TEV 3x3 study store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 08/29/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory